OpenStudy (anonymous):

A wire four feet long is cut in two pieces. One piece forms a circle of radius r, the other forms a square of side x. Choose r to minimize the sum of their areas. Then choose r to maximize. How do I do this problem and where do I even start? Very confused! :( Please explain from step by step? :/ Thanks!! :D

OpenStudy (jack1):

no one else is typing, so i'm gonna have a go, hope that's ok?

OpenStudy (anonymous):

yeah sure :)

OpenStudy (anonymous):

i honestly have no clue where to start haha

OpenStudy (jack1):

k, circumference of a circle = 2 pi r area of a circle = pi r^2 circumference of a square = 4s (s is side length) area of a square = s^2 now we know from question wording that: Circumference circle + circumference square = 4 ft we want: max area and min area ... with me so far @iheartfood ?

OpenStudy (anonymous):

yes:)

OpenStudy (jack1):

sweet so Circumference circle + circumference square = 4 ft 2 pi r + 4s = 48 inches now we want our answer in terms of r, so lets rearrange the above to eliminate s 4s = 48 - (2 pi r) s = 12 - 0.5 pi r still cool ? @iheartfood

OpenStudy (anonymous):

how did you find the 48 inches part?

OpenStudy (jack1):

1 foot = 12 inches so 4 ft = 48 inches

OpenStudy (anonymous):

ohh okay i see now :) i understand what you wrote above :)

OpenStudy (jack1):

cool cool, so now lets start playin with the area's Area circle + area square = lets call it z (z = total area) pi r^2 + s^2 = z and we know that: s = 12 - 0.5 pi r so \[\large pi \times r^2 + s^2 = z\] \[\large pi \times r^2 + (12 - 0.5 \times pi \times r)^2 = z\] \[\large pi \times r^2 + 144 - 0.25 \times pi^2 \times r^2 = z\] so to a few significant figures: \[\large 0.6742r^2+144 = z\] would you agree with that?

ganeshie8 (ganeshie8):

looks a "r" term is missing...

OpenStudy (anonymous):

should it be r^4 ? :/

OpenStudy (anonymous):

not quite sure hahaa

OpenStudy (jack1):

damn... so it is, sorry andd good catch @ganeshie8 !! ;) standby, my bad do over: \[\large pi \times r^2 + (12 - 0.5 \times pi \times r)^2 = z\] \[\large pi \times r^2 + 144 + 2.467r^2 - 37.7r = z \] \[\large 5.6086r^2 - 37.7r + 144 = z\] does that look a bit more realistic...?

OpenStudy (anonymous):

yes:) so from here, would we be using quadratic formula? not sure haha <--- just a guess on my part lol

OpenStudy (jack1):

so this equation, if u graph it, is still a parabola, and looks kinda like this: 

OpenStudy (anonymous):

haha yeah :P

OpenStudy (jack1):

quadratic formula would find where the parabola crosses the x axis in this case, we want to use derivatives to find where our lowest point is, as the above is a graph of z (area) as a function of r (radius)